1) You own a health club. The cost of fitness trainers is $1000 per month and $100 per class taught. Last month, 200 classes were taught.

a) Develop a cost equation for total cost of labor

b) What was the total variable labor cost last month ?

c) What was total labor cost last month ?

d) What was the unit cost of labor last month ?

2) Actual overhead is $1,000,000. Work in process is $100,000. Finished goods is $200,000. Cost of goods sold is $500,000.

Required:

1) Close out the overhead variance, assuming the amount of the variance is considered to be material.

2) Close out the overhead variance, assuming the amount of the variance is immaterial.
Answer #1

ANSWER

1)

a) Total cost = 1000+100 per class taught

b) Total variable cost = 100*200 = $20,000

c) Total labor cost = 20,000+1000 = $21,000

d) Unit cost = 21,000/200 = $105 per class taught

---------------------------------------------------------------------

DEAR STUDENT,
If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************